OpenStudy (anonymous):

360 gallons in 40 hours . the directions says write each rate or ratio in lowest terms change the same units when necessary

OpenStudy (anonymous):

In other words, how many gallons per hour ???

OpenStudy (anonymous):

i guess its my take home test its my make or break pass or fail :/

OpenStudy (anonymous):

\[\frac{ 360 g }{ 40 h } \to \frac{ 360 g }{ 40*3600s }\]

OpenStudy (anonymous):

it just says 360 gallons in 40 hours i know i have to change gallons into hours but idk how

OpenStudy (anonymous):

how do you change gallons to hours? i think they just want the rate to be reduced... your units will still be the same.

OpenStudy (anonymous):

I got 8 gallons an hours? is there a way to to check to see if its right

OpenStudy (anonymous):

360 gallons in 40 hours is the same as ???? gallons in 1 hour... find the ???.

OpenStudy (anonymous):

we cant count gallons in hours so we just have simplify the fraction

OpenStudy (anonymous):

8 gallons per hour

OpenStudy (anonymous):

check by multiplying: 8 * 40 = 360 ???? try again....

OpenStudy (anonymous):

its all im getting is 8

OpenStudy (anonymous):

\(\large \frac{360gallons}{40hours}=\frac{\cancel{360}^9gallons}{\cancel{40}^1hours}=\frac{9gallons}{1hour} \)